# When Do Transformers Learn Heuristics for Graph Connectivity?

### v1

@@ 5 Experiments
We test the theory in two parts. First, in §5.1 we verify the $3^{L}$ capacity threshold (Theorems 4.3 and 4.4) by measuring the maximum reliable path length one model can handle perfectly. Then, we trace training dynamics by projecting learned weights into the algorithmic $A\otimes I$ channel and the heuristic $B\otimes J$ channel (Theorem 4.6). Next, in §5.2 we introduce a simple data lever that up-weights within-capacity graphs, and shows this simple method suppresses the heuristic and promotes the algorithmic channel, as predicted by Theorems C.5 and C.9. Finally, we show this data lever prescribed by our theoretical analysis on disentangled transformers can transfer back to standard Transformers, and boost their generalization capabilities.
@@ 5.1 Capacity and Training Dynamics
$L$-layer Transformers Hit Their Capacity at Exactly $3^{L}$. We train disentangled transformers with 2 layers or 3 layers on Erdős-Rényi graphs with 24 or 64 nodes respectively. As shown in Figure 2, neither of the two models could make reliable predictions on node pairs $(u,v)$ with $d_{G}(u,v)>3^{L}$ but their predictions on node pair with $d_{G}(u,v)\leq 3^{L}$ are almost perfect with an $99\%$ accuracy. As shown in Figure 10, a 2-layer standard Transformer model also have the same capacity. It resonates with our exact capacity bound of disentangled transformers in Theorem 4.4 and justifies our dichotomy in Section 4.2. Overall, for any graph $G=(V,E)$, the decisive factor for transformer model depth is not simply the asymptotic $\Theta(\log|V|)$ relation to the number of nodes $n=|V|$ but more importantly the non-asymptotically exact relation to $\log_{3}\operatorname{diam}(G)$. It allows us to better understand the training dynamics in §4.3 with a dichotomy, something impossible without an exact relation. We note that our experiments did not enforce non-negativity of model weights and it demonstrates the predictive power of our theoretical analysis.
@@
Transformers Learn an Algorithm-Heuristic Mixture. In understanding the training dynamics, we first train a 1-layer disentangled transformer model on $\mathsf{ER}(n=8,p=0.2)$ graphs. We did not enforce any parameterization assumptions here. As show in Figure 3, a randomly initialized $W$ can converge to approximately a $A\otimes I_{n}+B\otimes J_{n}$ decomposition for some matrices $A,B\in\mathbb{R}^{2\times 2}$ and $J_{n}=\mathbf{1}\mathbf{1}^{\top}\in\mathbb{R}^{n\times n}$. Such parameterization also applies to deeper models as shown in Figure 9. They show the applicability of our decomposition Theorem 4.6. Then, we project the final weight $W$ onto this algebra as $\hat{W}=\hat{A}\otimes I_{n}+\hat{B}\otimes J_{n}$ by minimizing $\|W-\hat{W}\|_{F}$ and observe that the energy share (see §D) of $\hat{A}\otimes I_{n}$ in $\|W\|_{F}^{2}$ increases as training progresses but the share of $\hat{B}\otimes J_{n}$ first increases and then decreases, showing interesting training dynamics to be studied.
[…]
[width=]figures/restrict_diam/accuracy_comparison_exact_multiconfig.pdf
Figure 5: With 1-layer disentangled transformers with capacity $\mathsf{Cap}=3$ following Theorem 4.4, we vary the $d$ such that we restrict our training graphs to have $\operatorname{diam}(G)\leq d$. We also vary the edge probability of our training distribution $\mathsf{ER}(n=8,p=\cdot)$ for generality. We test on $\mathsf{2Chain}(n=8,k=\cdot)$ graphs with $k=2$ or 33 and show the exact match accuracy on configurations where the accuracy is non-zero for readability. We find if the training $d\leq\mathsf{Cap}$, models still learns the algorithmic solution up to problem size $d$ (see $d=2,k=2$ case on the left in orange) but fails to length generalize (see $d=2,k=3$ in orange on the right). On the other hand, if the training $d>\mathsf{Cap}$, model struggles to learn the algorithmic solution (see $d=4$ cases in red on both $k=2$ or 33). The best case overall is when setting $d=\mathsf{Cap}$, i.e., preventing the model from seeing beyond-capacity samples but still preserving at-capacity samples for better generalization. As shown in the green lines, with $d=3$, model achieves balanced testing accuracy on both $k=2$ and $3$.
[…]
[width=]figures/restrict_diam/beyond_capacity_proportion.pdf
@@
Figure 6: We vary the proportion of beyond-capacity graphs, and train the same disentangled transformer on stratified $\mathsf{ER}$ distribution and test on the same OOD $\mathsf{2Chain}$ distribution. We find that Transformers are robust towards a small amount of noises (beyond-capacity graphs). Although the $W$ is not exactly in the $A\otimes I_{n}$ form, the model still perform perfectly when the energy share of II-channel dominates (beyond roughly 90%). As shown on the right, there is a linear correspondence between model accuracy (averaged over individual node pairs) and algorithmic channel share.
5.2 Encouraging Transformers to Learn Algorithms over Heuristics
@@
Now that we understand the cause of why Transformers and disentangled transformer models learn heuristics that hurt their algorithmic computations (as shown in Figures 1 and 3), a natural question is whether we can mitigate it and encourage the models to up-weight the algorithm channel.
@@
Mitigation via the Data Lever. We propose a data-centric method: instead of training on all graphs from the $\mathsf{ER}$ distribution, we up-weight graphs whose $\operatorname{diam}(G)$ is within capacity following the dichotomy in Section 4.2. We dissect $\mathcal{G}=\mathcal{G}_{\leq}\sqcup\mathcal{G}_{>}$ into two sub-distributions where $\mathcal{G}_{\leq}=\{G\in\mathcal{G}:\operatorname{diam}(G)\leq 3^{L}\}$ only includes graphs containing no beyond capacity node pairs and $\mathcal{G}_{>}$ includes the rest. In Figure 4, we only train the 1-layer disentangled transformer on $\mathsf{ER}_{\leq}$, and then find the algorithmic $\hat{A}\otimes I_{n}$ channel is significantly promoted so that the learned weight only contains the algorithm channel. Furthermore, we find at-capacity graphs are crucial. In the case of Figures 5 and 11, where no graphs are to have $\operatorname{diam}(G)>2$, model also fails to learn generalizable solutions due to transformers’ poor length generalization abilities. It implies simply scaling up the model depth won’t equip it algorithmic capabilities naturally.
@@
Robustness towards Noise. To test the predictiveness of our theory, we evaluate if one beyond-capacity node pair is enough to encourage the model learning a heuristic-dominated method. We define $\rho(\mathcal{G})=\mathbb{E}_{G\in\mathcal{G}}{|\{(u,v)\in V,d_{G}(u,v)>3^{L}\}|}/{n^{2}}$ be the fraction of beyond-capacity node pairs in a graph distribution $\mathcal{G}$. In practice, $\rho$ can be controlled via stratified sampling from the mixture distribution $\mathcal{G}_{q}=q\mathcal{G}_{\leq}+(1-q)\mathcal{G}_{>}$. In Figure 6, we did stratified sampling between $\mathsf{ER}_{\leq}$ and $\mathsf{ER}_{>}$ and find that with a small $\rho(\mathcal{G})$, the model is still able to maintain high energy in the algorithm channel, and make perfect predictions on out-of-distribution $\mathsf{2Chain}$ graphs. It suggests that there exists a small $\rho^{\star}>0$ such that the model can still rely on the algorithm-channel to make predictions if the training distribution satisfies $\rho(\mathcal{G})\leq\rho^{\star}$.
[…]
[width=]figures/restrict_diam/roberta_n20_l2_plot.pdf
Figure 7: Standard transformer models learn generalizable solutions from within capacity data.
@@
Transferability to Standard Transformers models. Our theory from §4.3 makes a prescriptive suggestion to remove beyond capacity graphs to reduce dependence on heuristics, and Figure 4 demonstrated the effectiveness of this approach on disentangled transformers. We now evaluate this on standard Transformers. We train the same 2-layer Transformers model as in our preliminary study in §3.3 but this time, we train on the restricted distribution $\mathsf{ER}_{\leq}$ instead where all graphs $G\in\mathsf{ER}_{\leq}$ has $\operatorname{diam}(G)\leq 3^{2}=9$. As shown in Figure 7, when tested on the OOD $\mathsf{2Chain}$ dataset with maximum chain length 10, the one trained on $\mathsf{ER}_{\leq}$ can successfully generalize but the one trained on unconstrained distribution $\mathsf{ER}$ cannot.
@@ 6 Discussion and Conclusion
Acknowledgment
The authors acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing computing resources that have contributed to the research results reported within this publication. We also acknowledge the use of the USC NLP cluster provided by USC NLP Group. This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS250737 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. DF and RJ were also supported by gifts from the USC-Capital One Center for Responsible AI and Decision Making in Finance (CREDIF) and the USC-Amazon Center on Secure and Trusted Machine Learning. RJ was also supported by the National Science Foundation under Grant No. IIS-2403436. VS was supported by an NSF CAREER Award CCF-2239265, an Amazon Research Award, a Google Research Scholar Award and an Okawa Foundation Research Grant. The work was done in part while DF and VS were visiting the Simons Institute for the Theory of Computing. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the funding agencies.
@@ Appendix A Additional Details on Problem Setups and Preliminary Studies
Input and output. Given a simple graph on $n$ nodes with adjacency matrix $A\in\{0,1\}^{n\times n}$, let $\bar{A}=A+I_{n}$ be its self-loop augmented adjacency matrix. We treat $\bar{A}$ as input embedding: row $i$ is the token for node $i$; column $j$ indexes a feature tied to node $j$. The model outputs an $n\times n$ score matrix $\mathrm{TF}_{\Theta}^{L}(A)$, the predicted connectivity matrix.
@@
for $\ell=1,\ldots,L$ and heads $h=1,\ldots,H$. We use pre-norm residual blocks with LayerNorm ($\mathrm{LN}$) and GeLU activations. We do not use attention mask or any extra positional encoding; the identity in $\overline{A}$ already pins each token to a node.
@@
Metrics for Permutation Equivariance. Let $P\in S_{n}$ be the corresponding permutation matrix for any $\sigma\in\mathcal{S}_{n}$. For a given graph adjacancy matrix $A$, we compute the model’s prediction in respect to $P_{\sigma}$ as $\mathcal{M}(P_{\sigma}AP_{\sigma}^{\top})$. Now we define a equivariance consistency metric, Equivariance Consistency via Frobenius Cosine Similarity:
@@ B.2 Capacity
Assume the setup in Theorem 4.4. Suppose there exist some $n$-node graph, a layer index $\ell^{*}\in\{1,\ldots,L\}$, and vertices $u,v$ belonging to different connected components of the graph such that $(R_{\ell^{*}})_{u,v}>0$. Further assume that $\ell^{*}$ is globally minimal, in the sense that for all $n$-node graphs and all $\ell<\ell^{*}$, the corresponding $R_{\ell}$ has no false positive entries across components. Then there exists a graph $G$ such that $\mathrm{diam}(G)>3^{L}$, $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}>0$, where $u,v$ lie in different connected components of $G$.
@@
Step 2. Building a new graph. Initialize $G$ to the edgeless graph, keeping node isolated. We then embed $T_{u},T_{v}$ onto $G$ by adding edges according to the trees. Finally, we connect every vertex outside $S_{u}\cup S_{v}$ arbitrarily into a long chain.
We claim $G$ is the graph we seek. On one hand, every sum used by the certificates is a sum of nonnegative terms, and we have preserved a strictly positive summand at each step that appears in the tree. Hence $R_{\ell^{*}}(G)_{u,v}>0$ with $u,v$ also disconnected in $G$. On the other hand, under the choice of $n$ specified by Theorem 4.4, there exist at least $3^{L}+2$ vertices outside $S_{u}\cup S_{v}$, so connecting them into a long path guarantees $\mathrm{diam}(G)>3^{L}$. The claim then follows. ∎
@@
Assume the setup in Theorem 4.4. Further assume that for every $n$-node graph $G$ and every layer $\ell\in\{1,\ldots,L\}$, the post-ReLU scores $R_{\ell}(G)$ has no positive entry between distinct connected components of $G$. Then, for every graph $G$ and every $u,v\in V(G)$, if $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}>0$, we must have $\mathrm{dist}_{G}(u,v)\leq 3^{L}$. Consequently, if $G$ contains a connected component of diameter $3^{L}$ then $\mathsf{TF}_{\Theta}^{L}$ is not perfect on $G$.
@@
Step 2. Establishing the claim. We prove this claim via induction. The base case $\ell=0$ directly follows from the fact that $h_{0}(G)=[I_{n}\mid A(G)]$. For the inductive step, we assume that the claim holds at depth $\ell-1$ with radius $3^{\ell-1}$. As in Section B.2, there are two column types in $h_{\ell}$: inherited or newly appended columns. The former case is easy; if $c$ is inherited from $h_{\ell-1}$, then $h_{\ell}(G)_{i,c}=h_{\ell-1}(G)_{i,c}$, so the bound follows from the inductive hypothesis. We now assume $c$ is newly appended.
Suppose $(R_{\ell}(G)h_{\ell-1}(G))_{i,c}>0$ for a column $c$ with label $j$. Then there exists a row $k$ with $R_{\ell}(G)_{i,k}>0$ and $h_{\ell-1}(G)_{k,c}>0$. By the IH, $\mathrm{dist}_{G}(k,j)\leq 3^{\ell-1}$. Then we expand $R_{\ell}(G)_{i,k}>0$ to obtain column witnesses $p,q$, with $h_{\ell-1}(G)_{i,p}>0$, $h_{\ell-1}(G)_{k,q}>0$, and $(W_{\ell})_{p,q}>0$, as in Section B.2. Let $a,b$ be the labels of $p,q$, respectively. By IH again, $\mathrm{dist}_{G}(i,a)\leq 3^{\ell-1}$ and $\mathrm{dist}_{G}(k,b)\leq 3^{\ell-1}$. We now split the analysis into two cases.
  • If $a\neq b$, we derive a contradiction to the no-false-positives assumption by reusing the certificate procedure from Section B.2. Because $W_{\ell}\geq 0$ entrywise, every positive entry in $h_{t}(\bm{\cdot})$ admits a certificate supported on at most $s(t)\leq 2\cdot 3^{t}$ vertices. In particular, there exist certificates witnessing $h_{\ell-1}(G)_{i,p}>0$ (labeled $a$) and $h_{\ell-1}(G)_{k,q}>0$ (labeled $b$). Let $S_{a},S_{b}$ be the corresponding certificate vertex sets. Form a new graph $G^{\prime}$ on the same $n$ vertices whose connected components are two disjoint induced copies $S_{a}^{\prime},S_{b}^{\prime}$ of the subgraphs on $S_{a},S_{b}$ (leaving all other vertices outside $S_{a}\cup S_{b}$ isolated). Note this is feasible because $\lvert S_{a}\rvert+\lvert S_{b}\rvert\leq 4\cdot 3^{L-1}\leq n$ assumed by Theorem 4.4. By construction, there exist $i^{\prime}\in S_{a}^{\prime}$ and $k^{\prime}\in S_{b}^{\prime}$ with $h_{\ell-1}(G^{\prime})_{i^{\prime},p}>0$ and $h_{\ell-1}(G^{\prime})_{k^{\prime},q}>0$. Thus,
@@
The model’s output $\mathsf{TF}_{\Theta}^{L}(A(G))=h_{L}(G)W_{O}^{\top}$ is an entrywise nonnegative sum over the $n\times n$ blocks of $h_{L}(G)$. Since each block respects the $3^{L}$ locality bound, we have $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}=0$ whenever $\mathrm{dist}_{G}(u,v)>3^{L}$. Hence, on any graph whose largest component has diameter $3^{L}$, the model will inevitably miss a pair $(u,v)$ of nodes realizing this diameter. ∎

### v2

@@ 5 Experiments
We test the theory in two parts. First, in §5.1 we verify the $3^{L}$ capacity threshold (Theorems 4.3 and 4.5) by measuring the maximum reliable path length one model can handle perfectly. Then, we trace training dynamics by projecting learned weights onto the algorithmic $A\otimes I$ channel and the heuristic $B\otimes J$ channel (Theorem 4.7). Next, in §5.2 we introduce a simple data lever that up-weights within-capacity graphs, and show that this simple method suppresses the heuristic and promotes the algorithmic channel, as predicted by Theorems C.5 and C.9. Finally, we show this data lever prescribed by our theoretical analysis on Disentangled Transformers can transfer back to standard Transformers and boost their generalization capabilities.
@@ 5.1 Capacity and Training Dynamics
We train Disentangled Transformers with 2 layers or 3 layers on Erdős-Rényi graphs with 24 or 64 nodes respectively. As shown in Figure 2, neither of the two models could make reliable predictions on node pairs $(u,v)$ with $d_{G}(u,v)>3^{L}$ but their predictions on node pair with $d_{G}(u,v)\leq 3^{L}$ are almost perfect with an $99\%$ accuracy. As shown in Figure 9, a 2-layer standard Transformer model also has the same empirical capacity. These results resonate with our exact capacity bound of Disentangled Transformers in Theorem 4.5 and justifies our dichotomy in Definition 4.6. Overall, for any graph $G=(V,E)$, the decisive factor for Transformer model depth is not simply the asymptotic $\Theta(\log|V|)$ relation to the number of nodes $n=|V|$ but more importantly the non-asymptotically exact relation to $\log_{3}\operatorname{diam}(G)$. We also note that we do not enforce non-negativity of model weights during training, but our theoretical analysis remains predictive of model capacity.
Transformers Learn an Algorithm-Heuristic Mixture.
@@
To empirically understand training dynamics, we first train a 1-layer Disentangled Transformer model on $\mathsf{ER}(n=8,p=0.2)$ graphs, without enforcing any parameterization assumptions. As shown in Figure 3, a randomly initialized $W$ converges to a matrix approximately of the form $A\otimes I_{n}+B\otimes J_{n}$ for some matrices $A,B\in\mathbb{R}^{2\times 2}$. Deeper models also converge to such solutions, as shown in Figure 8. These results show the applicability of our decomposition Theorem 4.7. Then, we project the final weight $W$ onto this algebra as $\hat{W}=\hat{A}\otimes I_{n}+\hat{B}\otimes J_{n}$ by minimizing $\|W-\hat{W}\|_{F}$. We observe that the share (see §D) of $\hat{A}\otimes I_{n}$ in $\|W\|_{F}^{2}$ increases as training progresses but the share of $\hat{B}\otimes J_{n}$ first increases and then decreases. These provide empirical evidence supporting our two-phase story in §4.3.
Figure 4: Following insights from Theorems C.5 and C.9, we repeat the same experiment setup as in Figure 3 but only training on within-capacity graphs (see Definition 4.6). As shown in the solid lines, restricting training samples by capacity pushes the energy share of the algorithmic mechanism (the $A\otimes I_{n}$ channel) further to nearly 100% in the weight $W$. It simultaneously prevents the growth of the heuristic portion (the $B\otimes J_{n}$ channel).
@@
Figure 5: We vary the proportion of beyond-capacity graphs, and train the same Disentangled Transformer on stratified $\mathsf{ER}$ distribution and test on the same OOD $\mathsf{2Chain}$ distribution. We find that Transformers are robust towards a small amount of noises (beyond-capacity graphs). Although the $W$ is not exactly in the $A\otimes I_{n}$ form, the model still performs perfectly when the energy share of II-channel dominates (beyond roughly 90%).
5.2 Encouraging Transformers to Learn Algorithms
@@
Now that we understand why Transformers and Disentangled Transformer models learn heuristics that hurt their algorithmic computations (as shown in Figures 1 and 3), a natural question is whether we can mitigate this unwanted behavior and encourage the models to up-weight the algorithmic channel.
Mitigation via the Data Lever.
@@
We propose a data-centric method: instead of training on all graphs from the $\mathsf{ER}$ distribution, we up-weigh graphs whose $\operatorname{diam}(G)$ is within capacity following the dichotomy in Definition 4.6. We dissect the training distribution $\mathcal{G}$ into two sub-distributions: $\mathcal{G}_{\leq}=\{G\in\mathcal{G}:\operatorname{diam}(G)\leq 3^{L}\}$ only includes graphs containing no beyond capacity node pairs, and $\mathcal{G}_{>}$ includes the rest. In Figure 4, we only train the 1-layer Disentangled Transformer on $\mathsf{ER}_{\leq}$, and then find the algorithmic $\hat{A}\otimes I_{n}$ channel is significantly promoted so that the learned weight only contains the algorithm channel. Furthermore, we find at-capacity graphs are crucial. In the case of Figures 11 and 12, where no graphs are to have $\operatorname{diam}(G)>2$, the model also fails to learn generalizable solutions due to Transformers’ poor length generalization abilities. These results imply that simply scaling up the model depth does not naturally equip it with algorithmic capabilities .
Robustness to Noise.
@@
To test the predictiveness of our theory, we evaluate if a small amount of beyond-capacity node pairs is enough to encourage the model learning a heuristic-dominated method. We define $\rho(\mathcal{G})=\mathbb{E}_{G\in\mathcal{G}}{|\{(u,v)\in V,d_{G}(u,v)>3^{L}\}|}/{n^{2}}$ to be the fraction of beyond-capacity node pairs in a graph distribution $\mathcal{G}$. In practice, $\rho$ can be controlled via stratified sampling from the mixture distribution $\mathcal{G}_{q}=q\mathcal{G}_{\leq}+(1-q)\mathcal{G}_{>}$. In Figure 5, we performed stratified sampling between $\mathsf{ER}_{\leq}$ and $\mathsf{ER}_{>}$ and found that with a small $\rho(\mathcal{G})$, the model is still able to maintain high energy in the algorithm channel and make perfect predictions on out-of-distribution $\mathsf{2Chain}$ graphs. It suggests that there exists a small $\rho^{\star}>0$ such that the model can still rely on the algorithm-channel to make predictions if the training distribution satisfies $\rho(\mathcal{G})\leq\rho^{\star}$.
Figure 6: Standard Transformer models learn generalizable solutions from within capacity data. We train a 2-layer standard Transformer model on $\mathsf{ER}(n=20)$ graphs, with and without restricting graph diameters. When tested on OOD $\mathsf{2Chain}$ graphs, the one trained with the right data is able to generalize.
Transferability to Standard Transformers models.
@@
Our theory from §4.3 makes a prescriptive suggestion to remove beyond capacity graphs to reduce dependence on heuristics, and Figure 4 demonstrated the effectiveness of this approach on Disentangled Transformers. We now evaluate this on standard Transformers. We train the same 2-layer Transformers model as in our preliminary study in §3.3 but this time, we train on the restricted distribution $\mathsf{ER}_{\leq}$ instead where all graphs $G\in\mathsf{ER}_{\leq}$ have $\operatorname{diam}(G)\leq 3^{2}=9$. As shown in Figure 6, when tested on the OOD $\mathsf{2Chain}$ dataset with maximum chain length 10, the one trained on $\mathsf{ER}_{\leq}$ can successfully generalize but the one trained on unconstrained distribution $\mathsf{ER}$ cannot. It also helps model generalize to OOD $\mathsf{2Clique}$ graphs as well (see Figure 10).
@@ 6 Discussion and Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning, specifically our understanding of when neural networks learn generalizable algorithms versus brittle heuristics. Our findings have potential implications for improving the reliability and interpretability of machine learning systems deployed in safety-critical applications, where algorithmic correctness is paramount. By identifying data distribution properties that encourage algorithmic learning, this work may inform better training practices. We do not foresee direct negative societal consequences from this foundational research.
Appendix
@@ Appendix A Additional Details on Problem Setups and Preliminary Studies
Input and output. Given a simple graph on $n$ nodes with adjacency matrix $A\in\{0,1\}^{n\times n}$, let $\bar{A}=A+I_{n}$ be its self-loop augmented adjacency matrix. We treat $\bar{A}$ as input embedding: row $i$ is the token for node $i$; column $j$ indexes a feature tied to node $j$. To ease notation, we will simply write $A$ in place of $\bar{A}$ and always assume the adjacency matrix is self-loop augmented. The model outputs an $n\times n$ score matrix $\mathrm{TF}_{\Theta}^{L}(A)$, the predicted connectivity matrix.
@@
for $\ell=1,\ldots,L$ and heads $h=1,\ldots,H$. We use pre-norm residual blocks with LayerNorm ($\mathrm{LN}$) and GeLU activations. We do not use attention masks or any extra positional encoding; the identity in $\overline{A}$ already pins each token to a node.
@@
Metrics for Permutation Equivariance. Let $P\in S_{n}$ be the corresponding permutation matrix for any $\sigma\in\mathcal{S}_{n}$. For a given graph adjacency matrix $A$, we compute the model’s prediction in respect to $P_{\sigma}$ as $\mathcal{M}(P_{\sigma}AP_{\sigma}^{\top})$. Now we define a equivariance consistency metric, Equivariance Consistency via Frobenius Cosine Similarity:
@@ B.2 Capacity
Assume the setup in Theorem 4.5. Suppose there exist some $n$-node graph, a layer index $\ell^{*}\in\{1,\ldots,L\}$, and vertices $u,v$ belonging to different connected components of the graph such that $(R_{\ell^{*}})_{u,v}>0$. Further assume that $\ell^{*}$ is globally minimal, in the sense that for all $n$-node graphs and all $\ell<\ell^{*}$, the corresponding $R_{\ell}$ has no false positive entries across components. Then there exists a graph $G$ such that $\mathrm{diam}(G)=3^{L}+1$, $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}>0$, where $u,v$ lie in different connected components of $G$.
@@
Step 2. Building a new graph. Initialize $G$ to the edgeless graph, keeping node isolated. We then embed $T_{u},T_{v}$ onto $G$ by adding edges according to the trees. Finally, we connect $3^{L}+2$ vertices outside $S_{u}\cup S_{v}$ arbitrarily into a long chain of path length $3^{L}+1$. This is always possible because there are $n-|S_{u}\cup S_{v}|$ vertices outside the union, which is $\geq\big((7/3)\cdot 3^{L}+2\big)-4\cdot 3^{L-1}=3^{L}+2$, and this is why require $n\geq(7/3)\cdot 3^{L}+2$ in this Lemma.
We claim $G$ is the graph we seek. On one hand, every sum used by the certificates is a sum of nonnegative terms, and we have preserved a strictly positive summand at each step that appears in the tree. Hence $R_{\ell^{*}}(G)_{u,v}>0$ with $u,v$ also disconnected in $G$. On the other hand, under the choice of $n$ specified by Theorem 4.5, there exist at least $3^{L}+2$ vertices outside $S_{u}\cup S_{v}$, so connecting them into a long path guarantees $\mathrm{diam}(G)=3^{L}+1$. The claim then follows. ∎
@@
Assume the setup in Theorem 4.5. Further assume that for every $n$-node graph $G$ and every layer $\ell\in\{1,\ldots,L\}$, the post-ReLU scores $R_{\ell}(G)$ has no positive entry between distinct connected components of $G$. Then, for every graph $G$ and every $u,v\in V(G)$, if $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}>0$, we must have $\mathrm{dist}_{G}(u,v)\leq 3^{L}$. Consequently, if $G$ contains a connected component of diameter $3^{L}+1$ then $\mathsf{TF}_{\Theta}^{L}$ is not perfect on $G$.
@@
Step 2. Establishing the claim. We prove this claim via induction. The base case $\ell=0$ directly follows from the fact that $h_{0}(G)=[I_{n}\mid A(G)]$. For the inductive step, we assume that the claim holds at depth $\ell-1$ with radius $3^{\ell-1}$. As in Lemma B.1, there are two column types in $h_{\ell}$: inherited or newly appended columns. The former case is easy; if $c$ is inherited from $h_{\ell-1}$, then $h_{\ell}(G)_{i,c}=h_{\ell-1}(G)_{i,c}$, so the bound follows from the inductive hypothesis. We now assume $c$ is newly appended.
Suppose $(R_{\ell}(G)h_{\ell-1}(G))_{i,c}>0$ for a column $c$ with label $j$. Then there exists a row $k$ with $R_{\ell}(G)_{i,k}>0$ and $h_{\ell-1}(G)_{k,c}>0$. By the IH, $\mathrm{dist}_{G}(k,j)\leq 3^{\ell-1}$. Then we expand $R_{\ell}(G)_{i,k}>0$ to obtain column witnesses $p,q$, with $h_{\ell-1}(G)_{i,p}>0$, $h_{\ell-1}(G)_{k,q}>0$, and $(W_{\ell})_{p,q}>0$, as in Lemma B.1. Let $a,b$ be the labels of $p,q$, respectively. By IH again, $\mathrm{dist}_{G}(i,a)\leq 3^{\ell-1}$ and $\mathrm{dist}_{G}(k,b)\leq 3^{\ell-1}$. We now split the analysis into two cases.
  • If $a\neq b$, we derive a contradiction to the no-false-positives assumption by reusing the certificate procedure from Lemma B.1. Because $W_{\ell}\geq 0$ entrywise, every positive entry in $h_{t}(\bm{\cdot})$ admits a certificate supported on at most $s(t)\leq 2\cdot 3^{t}$ vertices. In particular, there exist certificates witnessing $h_{\ell-1}(G)_{i,p}>0$ (labeled $a$) and $h_{\ell-1}(G)_{k,q}>0$ (labeled $b$). Let $S_{a},S_{b}$ be the corresponding certificate vertex sets. Form a new graph $G^{\prime}$ on the same $n$ vertices whose connected components are two disjoint induced copies $S_{a}^{\prime},S_{b}^{\prime}$ of the subgraphs on $S_{a},S_{b}$ (leaving all other vertices outside $S_{a}\cup S_{b}$ isolated). Note this is feasible because $\lvert S_{a}\rvert+\lvert S_{b}\rvert\leq 4\cdot 3^{L-1}\leq n$ assumed by Theorem 4.5. By construction, there exist $i^{\prime}\in S_{a}^{\prime}$ and $k^{\prime}\in S_{b}^{\prime}$ with $h_{\ell-1}(G^{\prime})_{i^{\prime},p}>0$ and $h_{\ell-1}(G^{\prime})_{k^{\prime},q}>0$. Thus,
@@
The model’s output $\mathsf{TF}_{\Theta}^{L}(A(G))=h_{L}(G)W_{O}^{\top}$ is an entrywise nonnegative sum over the $n\times n$ blocks of $h_{L}(G)$. Since each block respects the $3^{L}$ locality bound, we have $\mathsf{TF}_{\Theta}^{L}(A(G))_{u,v}=0$ whenever $\mathrm{dist}_{G}(u,v)\geq 3^{L}+1$. Hence, on any graph whose largest component has diameter $3^{L}+1$, the model will inevitably miss a pair $(u,v)$ of nodes realizing this diameter. ∎